PLEASE SHOW WORK AND FORMULAS

What should be the prices of the following preferred stocks if comparable securities yield 7 percent? Why are the valuations different?
a. MN, Inc., $8 preferred ($100 par)
b. CH, Inc., $8 preferred ($100 par) with mandatory retirement after 20 years
Answer #1
a] Price = 8/7% = $      114.29
b] Price = 100/1.07^20+8*(1.07^20-1)/(0.07*1.07^20) = $      110.59
The valuations differ as,
the preferred stock at [a] is irredeemable paying
only interest perpetually, but
the preferred stock at [b] is redeemable paying
interest for 20 years and the par value at the end
of the 20th year.
Thus, since the cash flows differ the valuation also
differs.